# Spatio-Temporal Foundation Models: Vision, Challenges, and Opportunities

### v1

@@ III-A4 Scale Generalization
Spatio-temporal data spans a broad range of scales. Spatially, this can range from small-scale measurements, such as those at the meter level, to large-scale observations covering distances of hundreds of kilometers or more. Similarly, temporal scales vary from high-frequency observations, with timestamps in seconds or minutes, to low-frequency data collected over days or weeks. An STFM must be capable of generalizing across these diverse spatial and temporal scales.
@@ IV Current Research in STFMs
In this section, we examine the current landscape of STFM research, focusing on a selection of particularly recent and impactful models. We begin by briefly describing each model.
UniST [61] adopts a transformer-based encoder-decoder architecture, trained via masked-patch modeling where the objective is to reconstruct the original data from masked tokens in the model output. It uses various types of mask in pre-training to simulate several different problems in the reconstruction phase. It also constructs pools of learned prompts to encode different forms of spatio-temporal knowledge, e.g. spatial proximity and daily or weekly periodicity, which are added to the masked token embeddings to guide the model towards a better prediction. OpenCity [25] combines the transformer architecture with graph neural networks to learn spatio-temporal dependencies from large-scale, heterogenous traffic datasets. Alternatively, UrbanGPT [24] learns to encode ST sequences into new representations which can be incorporated into natural language prompts that can be understood and processed by a large language model. ClimaX [35] employs a vision transformer for various weather and climate-related tasks. It tokenizes and embeds each variable independently in order to flexibly handling varying numbers of input variables, and then aggregates across variables for a given region to reduce memory complexity. Pangu-Weather designs a 3D Earth Specific Transformer (3DEST) that processes weather information into cubic data and performs supervised training for various weather forecasting tasks.
IV-A Current STFM Generalization Capabilities
Table I shows our qualitative assessment of the generalization capabilities of the current state-of-the-art STFMs, based on their demonstrated performance in the original art. We now explain our assessment.
@@ IV-A Current STFM Generalization Capabilities
STFM Domain Spatial Temporal Scale
UniST [61]
UrbanGPT [24]
OpenCity [25]
Pangu-Weather [6]
ClimaX [35]
IV-B Domain Generalization
@@
An immediate observation is that, unlike FMs for other modalities, STFMs are highly fragmented by application. UniST, UrbanGPT and OpenCity are almost exclusively focused on traffic data Both UniST and OpenCity are trained and evaluated on 21 datasets in total, however the majority of these relate to traffic speed or flow in both cases, and a few other datasets for bike usage, taxi trajectories and cellular usage. Both models are evaluated in adapting to previously unseen applications by leaving certain datasets out of the training distribution entirely. On the other hand, UrbanGPT uses only four datasets from taxi, bike, and crime applications, three of which are used in pre-training.
Pangu-Weather and ClimaX are focused on a variety of atmospheric variables. ClimaX takes 48 input variables in total, but is only evaluated in forecasting 4 of them. Pangu-Weather, on the other hand, is only evaluated in predicting
IV-C Spatial Generalization
@@
Commonly used weather datasets use real observations from satellites with global coverage combined with numerical weather prediction models, which means they offer much broader spatial coverage compared with transportation datasets. In particular, ClimaX is trained on CMIP6 [32] and tested on ERA5 [43, 19], while Pangu-Weather is trained and evaluated on different years of data within ERA5. The primary experiments in both works test performance across the entire globe all at once, though ClimaX does also perform regional forecasting for North America only. However, as training and inference is performed over the entire globe at once, it does not demonstrate generalization from one seen regions to unseen regions. They also lack analysis of the distribution of errors in the spatial dimension.
IV-D Temporal Generalization
@@
Weather datasets generally contain much greater timespans, enabling training and evaluation over greater temporal ranges. Pangu-Weather is trained on 38 years of data from ERA5 (1979-2017), validated in 2019 and tested in 2018 and 2020-2021 data, while ClimaX is trained on CMIP6 which extends from 1850-2014, and tested on ERA5. Furthermore, it’s climate projection task extends until 2100, although it is important to note that this is not a temporal modelling task as it does not use historical observations to predict future states.
IV-E Scale Generalization
@@
Generalization across spatial scales is most relevant to the weather-related applications. Spatially, Pangu-Weather considers only one spatial resolution; $0.25\degree\times 0.25\degree$, which corresponds to approximately 28km $\times$ 28km cells. Comparatively, ClimaX evaluates performance with $5.625\degree\times 5.625\degree$ cells and $1.40625\degree\times 1.40625\degree$ cells. It also evaluates the model’s ability to perform climate down-scaling from 5.625°$\degree$ to 1.40625°$\degree$, as well as regional forecasting over North America only. However, these are all relatively large-scale and coarse-grained resolutions, which fails to capture the local patterns which are most useful for local-scale prediction.
Temporally, Pangu-Weather trains four separate models with different lead times (1, 3, 6 and 24 hours), and aggregates these models for any-time forecasting. This is in order to reduce the impact of error propagation in forecasting, especially at longer lead times, however it is also fundamentally at odds with the intended purpose of a foundation model. On the other hand, ClimaX randomizes lead times between 6 hours and 168 hours (1 week) in pre-training, and various other lead times for evaluation. Furthermore, they consider various forecasting tasks, including sub-seasonal prediction and climate projection with much greater timeframes.
@@ V Opportunities
V-A Cross-domain Synergies
Section III-A1 explored the complex relationships between ST sequences from different applications or sources. These relationships are severely under-utilized in existing ST models. For instance, based on our understanding of how infectious diseases spread through close contacts, we can infer that patterns of human mobility and traffic flow may significantly influence disease transmission. Therefore, a model that captures human mobility patterns could also provide valuable insights for modeling the spread of disease. By training STFMs to understand these patterns from data across multiple domains, we can enhance performance in specific applications by identifying and leveraging cross-domain correlations.
An important consideration is that, in many cases, relationships between applications are directional. For example, weather conditions likely impacts travel patterns, therefore accurate weather forecasting can aid in predicting traffic flow, but the reverse is not true. For this reason, a promising approach could be to design mechanisms to incorporate prior knowledge about spatio-temporal dynamics, such as physical laws and constraints, into the model. We can also look to recent advances in causal learning to uncover the nature of relationships between different features if their relationship is not already known. There is an established line of research dedicated to the discovery and inference of causal relationships in a spatio-temporal context [28, 13, 67] and this has continued in the era of spatio-temporal neural networks [66, 48, 52, 12].
V-B Unified Architectures
As discussed in Section II, ST data appears in a variety of different types and formats. An effective STFM should be able to handle all of these data types, however existing models are primarily or exclusively focused on only a single data type (namely, grid-based raster data). There have been some initial steps towards incorporating more flexible ST graph-based data very recently [62, 63], however this is mostly approached as a simple pre-processing step which may fail to account for the distinct properties of different data types. Furthermore, there is overwhelming focus on the transformer architecture, which suffers from quadratic complexity due in the self-attention mechanism. As the size of ST datasets and the number of variables and features to be modelled continues to grow, this complexity becomes a hindrance in the learning of ST patterns. Addressing these challenges requires innovative approaches, such as sparse attention mechanisms, efficient graph-based representations, and hybrid models that combine the strengths of transformers with other architectures like convolutional neural networks.
V-C Diverse Objectives
Existing ST research is primarily focused on predicting the spatial distribution of variables or forecasting their future states based on historical observations. Though these objectives have singificant practical importance, there are numerous other tasks that foundation models are capable of performing. Anomaly detection, for example, can be naturally derived from the prediction task by measuring the discrepancy between predicted and ground-truth sequences. In such cases, the quality of the underlying representation learning plays a crucial role in success, and a model that excels at one task is often also effective for others due to its robust feature learning. However, instilling STFMs with an ability to perform other tasks, such as classification, clustering, or recommendation, would greatly broaden their utility across various applications without compromising their accuracy in others. One effective approach is to add a lightweight module on top of the base foundation model, enabling it to leverage pre-trained features while adapting to new tasks in a flexible and efficient manner.
@@ V-D Adaptation to Distribution Shift
Foundation models are trained with large-scale data, across diverse sources and domains, which exposes them to a wide range of scenarios. Nevertheless, during inference, they remain heavily reliant on the statistical properties of their training data. If the test data deviates from the training distribution (e.g., due to temporal changes like urban development or spatial differences like regional policies), the model’s performance can drop sharply. Moreover, the particularly high complexity and dimensionality of spatio-temporal data increases the likelihood of encountering unseen patterns or out-of-distribution (OOD) scenarios in practice. This challenge is identified in Sections III-A2 and III-A3. Foundation modal adaptation offers a promising approach to address this challenge. There has been a great amount of research interest in the adaptation of foundation models for computer vision to boost their performance in specific tasks involving data outside their original training distribution. Recent research focuses on techniques like domain-adversarial training, which encourages the model to learn domain-invariant features, and meta-learning to rapidly adapt to new distributions with limited data from new distributions.

### v2

@@ III-A4 Scale Generalization
Spatio-temporal data spans a broad range of scales. Spatially, this can range from small-scale measurements, such as those at the meter level, to large-scale observations covering distances of hundreds of kilometers or more. Similarly, temporal scales vary from high-frequency observations, with timestamps in seconds or minutes, to low-frequency data collected over days, weeks, or months. STFMs must be capable of generalizing across these diverse spatial and temporal scales.
@@ IV Current Research in STFMs
In this section, we explore the current state of research on STFMs. As of this writing, we found no studies that comprehensively address the full range of spatial-temporal applications outlined earlier. Nevertheless, we identified six STFMs that align with the principles of foundation models as defined in Definition III.2, three of which primarily focus on transportation-related applications and the other three focus on weather-related applications. We now provide a brief overview of these models.
@@
STFM Data Type Domain Spatial Temporal Scale
UniST [77] Grid Traffic speed, cellular usage, taxi demand, bike usage, crowd flow [Uncaptioned image] Various US and Chinese cities [Uncaptioned image] 2013-2022 [Uncaptioned image] [Uncaptioned image]
UrbanGPT [33] Time-series Taxi usage, bike usage, crime [Uncaptioned image] New York City, Chicago [Uncaptioned image] 2017, 2021 [Uncaptioned image] [Uncaptioned image]
OpenCity [34] Grid, sensor Traffic flow, taxi demand, bike trajectory, traffic speed [Uncaptioned image] Various US and Chinese cities [Uncaptioned image] 2013-2022 [Uncaptioned image] [Uncaptioned image]
W-MAE [42] Grid ERA5 reanalysis data [Uncaptioned image] Global coverage [Uncaptioned image] 1979-2018 [Uncaptioned image] [Uncaptioned image]
Pangu-Weather [6] Grid ERA5 reanalysis data [Uncaptioned image] Global coverage [Uncaptioned image] 1979-2021 [Uncaptioned image] [Uncaptioned image]
ClimaX [47] Grid CMIP6 simulations, ERA5 reanalysis data [Uncaptioned image] Global coverage [Uncaptioned image] 1850-2014 [Uncaptioned image] [Uncaptioned image]
UniST [77], OpenCity [34], and UrbanGPT [33] are three STFMs designed for transportation applications. In particular, traffic prediction is the primary focus of both UniST and OpenCity, both of which adopt the transformer architecture to learn from large-scale, heterogeneous traffic datasets. UniST adopts a masked-patch modelling pre-training objective, where grid-based ST data is organised into ST patches, and the model is trained to reconstruct the data of masked patches given the unmasked patches. Various types of mask are used in pre-training to simulate different problem settings, and it also learns prompt memory pools to encode different forms of spatio-temporal knowledge for improved reconstruction. OpenCity instead combines a transformer with a graph neural network to learn distinct spatial and temporal patterns through separate attention layers and is trained with an absolute error loss function. On the other hand, UrbanGPT trains a spatio-temporal dependency encoder to a sequence of ST embedding tokens with the same dimensionality as the internal representation space of a large language model (LLM). The LLM is then used as the processing engine, given contextual information and a task description in the form of a natural language instruction, to predict the next tokens in the sequence, which are converted back into ST data through a regression layer.
ClimaX [47], Pangu-Weather [6], and W-MAE [42] are three weather-focused STFMs. In this context, there are two massive-scale broad datasets for weather science: CMIP6 climate simulations and [44], and ERA5 reanalysis data [57, 24]. Both of these datasets offer global coverage of a wide variety of atmospheric variables over extremely long temporal horizons. The primary focus of these STFMs is the ability to forecast the future state of several of these variables simultaneously. ClimaX is based on a vision transformer, and it tokenizes each variable independently in order to flexibly handle different sets of input variables. To reduce memory complexity, it aggregates these embeddings to a single vector representation for each spatial unit. The model is trained to minimize the latitude-weighted prediction error. Pangu-Weather designs a 3D Earth Specific Transformer (3DEST) that merges surface-level and upper-air variables into cubic data patches, and employs a encoder-decoder architecture to learn their patterns, also trained with an error-based loss function. On the other hand, W-MAE employs a masked auto-encoder pre-training objective with a ViT backbone architecture for multi-variable weather forecasting. It is important to note that all three weather foundation models use a single time-stamp of historical observations as input to predict the future state of the atmosphere at a variety of lead times, which means they do not explicitly model temporal dependencies in order to make predictions.
Table I shows our qualitative assessment of the generalization capabilities of these STFMs, based on their demonstrated performance in their original work. We considered two main factors when determining our assessment: (1) does the choice of data used in pre-training contribute towards this form of generalization? and (2) is there an effort to experimentally evaluate or demonstrate this form of generalization capability? We denote a positive assessment (i.e. significant evidence) as [Uncaptioned image], a neutral assessment (i.e. some but insufficient evidence) as [Uncaptioned image], and a negative assessment (i.e. little to no evidence) as [Uncaptioned image]. We emphasize that our assessment should not be interpreted as a reflection or comparison of their performance, rather a study of the extent to which current research aligns with the principles of foundation models.
IV-1 Domain Generalization
@@
As mentioned earlier, our immediate observation is the significant fragmentation between high-level applications. UniST, UrbanGPT and OpenCity are exclusively focused on transportation applications. Both UniST and OpenCity are both trained or evaluated on 21 datasets in total, but the majority of these relate to traffic speed and traffic flow, with only a few other datasets for other applications like bike usage, taxi demand and cellular usage. Both models are evaluated in both ID generalization by training on multiple applications simultaneously, as well as OOD generalization by leaving certain datasets out of the training distribution. On the other hand, UrbanGPT uses only four datasets: taxi, bike, and crime datasets from New York City and a taxi dataset from Chicago. The model is tested in ID generalization across the three training datasets, but not OOD generalization.
As for weather-focused STFMS, there are a greater range of features in the form of different atmospheric variables, however all three models evaluate perdiction performance for only 4 of them. Pangu-Weather is also evaluated with a tropical cyclone tracking task, which is derived from the local minmum of mean sea level pressure.
IV-2 Spatial Generalization
@@
As mentioned, the two weather datasets which are used in both pre-training and evaluating the weather foundation models are global in their spatial coverage. ClimaX is trained on CMIP6 [44] and fine-tuned and tested on ERA5 [57, 24], while Pangu-Weather and W-MAE are trained and evaluated on different years of data from ERA5. Evaluation of all models is performed across the entire globe in aggregate, though ClimaX also performs regional forecasting for the North America region specifically. As both training and inference is performed over the entire globe, this does not demonstrate generalization from one seen regions to unseen regions, and also lacks fine-grained analysis of how errors are spatially distributed across different regions.
IV-3 Temporal Generalization
@@
The weather datasets generally contain much greater timespans, enabling training and evaluation over greater temporal ranges. ClimaX is trained on data from 1850 to 2014 from CMIP6, and tested on ERA5 from 1979 to 2018. Pangu-Weather is trained on 38 years of data from ERA5 (1979-2017), validated in 2019 and tested in 2018 and 2020-2021 data, while W-MAE trains two separate models: one with 2 years and one with 37 years of data.
IV-4 Scale Generalization
@@
Generalization across spatial scales is most relevant to the weather-related applications. Spatially, Pangu-Weather considers only one spatial resolution; $0.25\degree\times 0.25\degree$, which corresponds to approximately 28km $\times$ 28km cells. Comparatively, ClimaX evaluates performance with $5.625\degree\times 5.625\degree$ cells and $1.40625\degree\times 1.40625\degree$ cells. It also evaluates the model’s ability to perform climate down-scaling from 5.625°$\degree$ to 1.40625°$\degree$, as well as regional forecasting over North America only. However, these are all relatively large-scale and coarse-grained resolutions, which fails to capture the local patterns which are most useful for local-scale prediction. Temporally, Pangu-Weather trains four separate models with different lead times (1, 3, 6 and 24 hours), and aggregates these models for any-time forecasting. This is in order to reduce the impact of error propagation in forecasting, especially at longer lead times, however it is also fundamentally at odds with the intended purpose of a foundation model. On the other hand, ClimaX randomizes lead times between 6 hours and 168 hours (1 week) in pre-training, and various other lead times for evaluation. Furthermore, they consider various forecasting tasks, including sub-seasonal prediction and climate projection with much longer temporal horizons.
Overview: We make the following general observations from our assessment of existing STFM research.
  • Existing STFMs are fragmented into transportation foundation models and weather foundation models with no overlap in applications or evaluations. Within each of these categories, the transportation foundation models are primarily focused on the task of traffic prediction, with a few datasets from applications like taxi demand, bike usage and crowd levels either included in the training sets or used to measure OOD generalization in the test set. On the other hand, weather foundation models often incorporiate a wide variety of atmospheric variables in their input variables, but are mostly evaluated in their prediction of only a small subset of these variables during inference.
  • There are numerous datasets used to train and evaluate transportation foundation models, however they are critically limited in spatial coverage to just a small selection of major urban centers, almost entirely limited to cities located in the USA and China.
  • Weather foundation models are exclusively trained on one or both of two massive-scale weather datasets, CMIP6 and ERA5, which offer global spatial coverage over extremely long temporal horizons. However, they are both limited to coarse granularities which limits their ability to generalize to applications which require more fine-grained analysis over smaller spatial regions.
@@ V Opportunities
In this section, we outline a range of potential avenues and directions for future research, both to address the limitations identified in the previous section, and also to unlock new performance enhancements and capabilities.
V-A Unified Architectures
As discussed in Section II, ST data exists in a variety of types and formats. However, existing STFMs predominantly focus on rasterized grid-based data, often requiring data in other formats to be transformed into a structured grid, which can introduce significant information loss and obscure fine-grained spatiotemporal relationships. Ideally, STFMs should be designed to handle heterogeneous ST data natively, rather than relying on lossy pre-processing techniques. A number of recent works have started to explore the integration of more flexible ST graph-based data representations [78, 79]. However, these approaches are still in their early stages and often treat graph-based structures as a secondary input, rather than a primary modeling component. This can lead to suboptimal representations that fail to fully capture the complex spatial dependencies and topological relationships inherent in many ST datasets. A more principled approach is needed to develop STFM architectures that seamlessly integrate structured and unstructured data.
Another limitation is the reliance on transformer-based architectures. While transformers have demonstrated strong performance in many ST applications, their quadratic complexity due to the self-attention mechanism presents a significant scalability challenge, particularly as ST datasets grow in size and complexity. This issue is exacerbated when dealing with high-resolution spatial data, long temporal sequences, or multi-modal inputs that require modeling diverse interactions across time and space. Addressing these challenges necessitates novel architectural innovations, such as sparse attention mechanisms, hierarchical modeling techniques, and hybrid approaches that combine the strengths of transformers with more computationally efficient architectures like convolutional neural networks (CNNs), recurrent neural networks (RNNs), and graph neural networks (GNNs).
Furthermore, there is a pressing need for more standardized benchmarks and evaluation protocols for STFMs. Existing studies use disparate datasets and evaluation settings. For example, UniST evaluates short-term and long-term prediction performance using a temporal horizon of 6 time-steps and 64 time-steps respectively, whereas OpenCity uses 1 day of observations, which means the number of time-steps varies between datasets depending on their frequency. Similarly, inconsistencies in spatial resolutions, training and test applications, evaluation metrics, and ID vs OOD generalization settings make it challenging to compare performance between models. Establishing comprehensive ST benchmarks that encompass diverse data types and downstream tasks would provide a clearer picture of model capabilities and drive further advancements.
V-B Cross-domain Synergies
Section III-A1 explored the complex, under-utilized relationships between spatio-temporal (ST) sequences from different applications or sources. Current ST models often treat each application in isolation, neglecting the rich information shared across domains. For example, based on our understanding of how infectious diseases spread through close contact, we can infer that patterns of human mobility are intricately linked to the spread of transmissible diseases within a population. For this reason, a model that accurately captures traffic flow and population movement can offer valuable insights for predicting disease outbreaks, informing public health interventions, and ultimately mitigating their impact. By training STFMs to recognize and leverage these cross-domain correlations, we can unlock significant advancements in individual applications. A model trained on diverse datasets, encompassing weather patterns, traffic data, social media activity, and epidemiological information, could learn more robust and generalizable representations of spatio-temporal dynamics, leading to improved performance in each individual domain.
Crucially, the relationships between different domains are often directional. Weather patterns influence travel behavior, making accurate weather forecasting a valuable input for predicting traffic flow. However, the reverse is not true; traffic congestion does not significantly impact the weather. Therefore, incorporating prior knowledge about these directional relationships, such as physical laws and constraints, is essential for building effective STFMs. For instance, incorporating known relationships between temperature, humidity, and evapotranspiration could enhance agricultural yield prediction models.
When the nature of the relationships between features is unknown, recent advances in causal learning offer powerful tools for discovery. Causal inference techniques can help disentangle correlation from causation, revealing the true drivers of spatio-temporal phenomena. This is particularly important for understanding complex systems where multiple factors interact in intricate ways. While established research exists on discovering and inferring causal relationships in spatio-temporal data [39, 16, 83], the rise of spatio-temporal neural networks has opened new avenues for exploration. Recent work has begun to integrate causal reasoning into these models, including generative approaches [82], graph neural networks for causal discovery [63], methods for deciphering causal relationships from observational data [68], and frameworks for building causally aware ST models [15]. Future research should focus on developing novel methods for incorporating causal knowledge directly into STFMs, enabling them to learn more interpretable, robust, and generalizable representations of the world. This includes exploring how to represent causal relationships within the model architecture, how to train models to respect causal constraints, and how to leverage causal information for improved prediction and forecasting. Furthermore, investigating how to handle noisy and incomplete data in causal discovery within a spatio-temporal context is crucial for real-world applications.
V-C Multi-modal Training
Information in the physical world is inherently multi-modal, encompassing diverse data sources such as numerical time-series, remote sensing imagery, textual descriptions, audio signals, and more. Integrating multi-modal learning into spatio-temporal foundation models (STFMs) has the potential to significantly enhance their ability to capture complex dependencies and improve their robustness across real-world challenges. By leveraging complementary information from different modalities, STFMs can achieve a more holistic understanding of spatio-temporal phenomena, reducing uncertainty and improving predictive performance.
For example, predicting crop yields could benefit from fusing satellite imagery (providing information on vegetation health and land cover), weather time-series data (temperature, rainfall), and soil sensor readings (moisture, nutrient levels). Similarly, for urban traffic forecasting, combining traffic camera footage (visual flow), GPS data from vehicles (speed and location), and social media posts (reports of accidents or road closures) could lead to more accurate and robust predictions. In disaster response, multi-modal data like satellite imagery (damage assessment), weather radar (storm tracking), and social media reports (citizen accounts of the situation) can provide a more comprehensive understanding of the event and improve the efficiency of relief efforts. Maps serve as a particularly crucial source of structured knowledge and constraints for learning from spatial-temporal data. As a structured representation of the contextual characteristics of the environment, maps can help to provide hard constraints, such as physical barriers that restrict movement (e.g., rivers, buildings, or restricted zones), and soft constraints, such as land-use characteristics that influence event likelihoods (e.g., residential versus commercial areas). They also serve as a regularization mechanism, guiding model learning through spatial smoothing techniques and topology-aware constraints, such as enforcing connectivity in road network-based predictions using graph neural networks.
There have recently been some initial strides towards curating large-scale, multi-modal datasets for spatio-temporal applications. For example, the Terra dataset [14] encompasses 45 years’ worth of hourly time-series data from over 6 million grid areas worldwide, covering various meteorological variables while incorporating multi-modal information such as explanatory text, geo-imaging, and satellite imagery. However, these efforts remain scarce and largely limited to a few applications and modalities.
One major challenge in multi-modal ST learning is the alignment and fusion of heterogeneous data sources, which often have vastly different temporal resolutions, spatial granularities, and noise characteristics. For instance, while satellite imagery may be available at a high spatial resolution but captured at irregular intervals, time-series sensor data may provide continuous readings but at a coarser spatial resolution. Effective STFMs must develop novel alignment techniques to efficiently integrate such disparate data sources without introducing artifacts or losing critical information.
@@ V-D Adaptation to Distribution Shift
Foundation models are trained with large-scale data, across diverse sources and domains, which exposes them to a wide range of data distributions. Nevertheless, during inference, they remain heavily reliant on the statistical properties of their training data. If the test data significantly deviates from the training distribution, the model’s performance can drop sharply. In practice, the physical environment is constantly evolving which guarantees some form of distribution shift in data over time, e.g. temporal changes like urban development or changes in regional policies. Moreover, the particularly high complexity and dimensionality of spatio-temporal data increases the likelihood of encountering unseen patterns or out-of-distribution (OOD) scenarios in practice.
Foundation model adaptation [38, 11, 29] offers a promising approach to address this challenge and has received significant research attention in order to boost their performance in individual vision and language tasks, particularly involving data outside the original training distribution. Several adaptation strategies have shown promise for addressing OOD challenges. For example, domain-adversarial training encourages the model to learn representations invariant to the specific domain of the training data. By minimizing the discrepancy between feature distributions of different domains, the model becomes more robust to domain shifts and generalizes better to unseen data. In the ST context, this could involve training the model to be invariant to specific geographic regions or time periods, enhancing adaptability to new locations or future conditions. Meta-learning aims to train models that can rapidly adapt to new tasks or distributions with only a few examples. The model learns a set of parameters well-suited for fast adaptation, enabling quick adjustments to new ST patterns with limited data. This is particularly relevant for ST applications where labeled data in new domains might be scarce. Test-time adaptation adjusts the model’s parameters during inference using the unlabeled test data itself, proving effective for handling gradual distribution shifts over time. Techniques like self-training and entropy minimization can be used to adapt the model to the characteristics of the current test batch, improving performance on unseen data.